A coin collector has 61 dimes and nickels with a total face value of $5.25. How many of each coin does she have?
Answer by umnikmo(75)

X-dimes
Y-Nickels
X+Y=61
10X+5Y=525
X=61-Y
10(61-Y)+5Y=525
610-10Y+5Y=525
85=5Y
Y=17
X=61-17=44.
CHECK:
10*44+5*17=525